OpenStudy (anonymous):

in 1996, the number of kidney transplant procedures K was approximately 12264. From 1996 to 2002, the number of kidney transplant procedues increased by an average of 586 procedures per year. I need to write a verbal model that gives the number of kidney transplant procedures t years after 1996.

OpenStudy (anonymous):

t=1,2,3,4,5

OpenStudy (anonymous):

i can't quite grasp what you are trying to ask

OpenStudy (anonymous):

12264+(586t)

OpenStudy (anonymous):

need an algebraic model that gives the number of kidney transplants procedures t years after 1996

OpenStudy (anonymous):

just need to know how to go about figuring this out

OpenStudy (anonymous):

well 586t wil give you how much each year it's increasing by

OpenStudy (anonymous):

so it doesn't matter that 586 is the average?

OpenStudy (anonymous):

The question is asking for an approximation, there is no way to accurately predict the future.

OpenStudy (anonymous):

Okay, so the first year, t=0, there were 12264 transplants. Every year after, the number of procedures increased by approximately 586 procedures. So to find the approximate number of procedures the next year, t=1, you take the number of procedures the first year, y(t=0) and add 586. y(t=1) = y(t=0) + 586. Then you do the same thing again. The number of transplants two years later, y(t=2) is y(t=1) + 586. Now you can start to see a pattern. At any given time, lets call it n+1, you can find it using t=n. y(t=n+1) = y(t=n) + 586. Since you know the exact value, you can approximate for any future t.

OpenStudy (anonymous):

That is how Outkast3r09 came to her equation. They started with the fixed number, 12264, and knowing that each year the number goes up by approximately 586, they added 586 for each year after 1996.